Column 2 Row35A49.3 159.26 463.54 161.02 483.55 124.84 575.85 147.71 567.66 193.22 567.81 65.2 306.68 616.68 612.56 720.15 460.03 782.62 375.67 728.84 424.35 979.52 360.48 977.33103.9 53.8 56.9 60.6 4 6 7 8 9 10 72.6 82.4 90.9 96.5 99.6 3 OF 4

6. Calculate the mean and variance of Xij

7. Calculate the mean and variance of Xi2-
Answer #1

(6)

(i)

Mean = ar{x} = 3325.82/11 = 302.3473

x (x - ar{x} ) (x - ar{x} )2
159.26 - 143.0873 20473.9676
161.02 - 141.3273 19973.3980
124.84 - 177.5073 31508.8319
147.71 - 154.6373 23912.6861
193.22 - 109.1273 11908.7617
306.68 4.3327 18.7725
612.56 310.2127 96231.9362
460.03 157.6827 24863.8425
375.67 73.3227 5376.2223
424.35 122.0027 14884.6655
360.48 58.1327 3379.4140
Total = 252532.4982

Variance is given by:

Σ2l (ri-i)2-252532. 4982- 25253.2498 10 10

So,

Answer is:

25253.2498

(7)

(i)

Mean = ar{x} = 7463.55/11 = 628.5045

x (x - ar{x} ) (x - ar{x} )2
463.54 - 214.9645 46209.7558
483.55 - 194.9545 38007.2748
575.85 - 102.6545 10537.9557
567.66 - 110.8445 12286.5133
567.81 - 110.6945 12253.2824
616.68 - 61.8245 3822.2744
720.15 41.6455 1734.3439
782.62 104.1155 10840.0279
728.84 50.3355 2533.6580
979.52 3301.0155 90610.3039
977.33 298.8255 89296.6523
Total = 318132.0423

Variance is given by:

(r-r)2 318132042331813.2042 10 10

So,

Answer is:

31813.2042